OpenStudy (anonymous):

Find the complex square roots and write the answer in polar form. 4(cos60º + i sin60º)

Parth (parthkohli):

Again, convert this complex number to the form \(\large re^{i\theta}\). Possibly any square root you will encounter is just when you raise something to the power \(1/2\). So you should first get the exponential form, then raise that to the power \(1/2\).

Parth (parthkohli):

Just another mention:\[\large \cos(\theta) + i\sin(\theta) = e^{i\theta}\]

OpenStudy (gudden):

@Dddd1123 is it clear ?